OpenStudy (anonymous):

Find an expression for the general term of the series below. Assume the starting value of the index, k, is 0.

OpenStudy (anonymous):

\[(x-5)^4 - \frac{ (x-5)^7 }{ 2!} + \frac{ (x-5)^{10} }{ 4! } - \frac{ (x-5)^{13} }{ 6! }+ . . .\]

OpenStudy (anonymous):

@satellite73 can you help me please

OpenStudy (raden):

the odd terms always negative, while even terms positive... it means the terms has (-1)^(n+1), for n is natural number

OpenStudy (raden):

the power of (x-5) is the aritmetic sequence with the terms 4,7,10,13,...

OpenStudy (raden):

just use the formula : an = a1 + (n-1)d an = 4 + (n-1)3 an = 3n + 1

OpenStudy (raden):

for the factorial's 0!, 2!, 4!, 6!, ... so on use the arithmetic sequence too

OpenStudy (raden):